1. A student who was in a hurry to complete this experiment did not follow the exact directions for some of the procedural steps. Briefly explain what effect, if any, the following changes would have on the whole number ratio of Mg and O in the compound the student produced.

(1) When arriving at the laboratory station, the student found a dirty crucible and cover. In the interest of time, the student skipped steps 2-8, weighed the cruible and cover, and proceeded directly to Step 9

(2) While heating the Mg in the crucible, the cover fell onto the laboratory bench and broke. The student found a new cover and used it for the remainder of the experiment

(3) After adding water to convert and Mg3N2 to magnesium oxide, the student completely covered the crucible while doing the final heating.

(4) Because the cruicble seemed to be taking forever to cool, the student decided to make the final weighing when the crucible was still slighty warm

2. A sample of a compound containing C,O and silver (Ag) weighed 1.372 g. On analysis it was found to contain .288g O and .974 g Ag. The molar mass of the compund is 303.8 g mol and the molar masses of C, O and Ag are 12.01 , 16.00 g mol, and 107.9 g mol, respectively. Determine the empirical and molecular formulas of the compound

3. Briefly explain how your Mg to O ratio would have been affected if you had failed to convert all of the Mg3N2 present to magnesium oxide in Steps 18-22 of the procedure
Answer #1

1. If crucible is not washed and dried properly, the dirt in the crucible will increase the weight of the crucible. After the experiment the weight of the crucible may not be same if some dirt adhered to the outside of the crucible falls of or removed while handling it during the experiment.

If that happens total mass of the oxide calculated after the experiment will come to be larger than what it actually is. Thus the ratio of Mg and O will appear to be smaller than reality.

Also, if the crucible contained some other compounds initially which reacts with the gas inside the crucible on heating, then total mass of the oxide calculated after the experiment will appear to be bigger than what is is actually and the ratio of Mg and O will appear to be smaller than reality.

2. If the new lid is heavier than the original one then total mass of the oxide calculated after the experiment will appear to be bigger than what it is actually and the ratio of Mg and O will appear to be smaller than reality.

If the new lid is lighter than the original one then total mass of the oxide calculated after the experiment will appear to be smaller than what is actually and the ratio of Mg and O will appear to be larger than reality.

3. When water is added to convert Mg3N2, NH3 vapor and water vapor produces in the crucible. If the crucible is completely covered then the ammonia and water vapor will remain in the crucible and total mass of the oxide after the experiment will appear to be larger than reality.

So the ratio of Mg and O will appear to be smaller than what it actually is.

4. If the crucible is warm, the metal part of the balance may expand which will ultimately contribute to an error. Moreover warm interior of the crucible can create an air current inside and around the crucible which will give an unsteady and incorrect result.

2)

Mass of sample= 1.372gm, mass of sample = mass of C+ mass of O + mass of Ag= 1.372 gm

Mass of C+ 0.288+0.974=1.372 gm=0.11 gm

Moles = mass/molar mass

Moles : C= 0.11/12.01= 0.0092, O=0.288/16=0.018, Ag= 0.974/107.9 =0.0090

Molar ratio of C: O:Ag= 0.0092: 0.018:0.0090= 0.0092/0.009: 0.018/0.009: 0.0090/0.0090= 1:2:1

So the empirical formula is CO2Ag

Molar mass of empirical formula is = 12+32+107.9= 151.9

Molar mass of empirical formula*n= molar mass , where n= no of empirical formulas

Hence n*151.9= 303.8, n= 2. Hence Molecular formula= C2O4Ag2